# A geometric approach toconjugation-invariant random permutations

### v3

@@ 7.1 Dependency graphs and asymptotic normality
Let $t$ be a cycle type of size $n$, and $\mathcal{L}_{t}$ be the graph defined in Section 1.3. The following lemma is a direct consequence of the definitions.
@@
In the geometric construction of 1.1, the graph $\mathcal{L}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$ is a dependency graph for the variables $(Z_{i})_{i\in\mathcal{V}_{t}}$.
Now we can use $\mathcal{L}_{t}$ to construct another dependency graph. Let $\binom{\mathcal{V}_{t}}{r}$ denote the set of subsets of size $r$ in $\mathcal{V}_{t}$. Define a graph $\binom{\mathcal{L}_{t}}{r}$ with vertices $\binom{\mathcal{V}_{t}}{r}$ as follows: $I,J\in\binom{\mathcal{V}_{t}}{r}$ share an edge if and only if there exist $i\in I$ and $j\in J$ such that $i=j$ or $i\sim j$ in $\mathcal{L}_{t}$. Using 7.1, it is straightforward to check that $\binom{\mathcal{L}_{t}}{r}$ is a dependency graph for the family of random variables (.$\left(\big{.}(Z_{i})_{i\in I}\right)_{I\in\binom{\mathcal{V}_{t}}{r}}$.
The idea that the variables $(Z_{i})_{i\in\mathcal{V}_{t}}$ are “partly dependent” is motivated by the fact that these graphs have low degrees: the degrees of $\mathcal{L}_{t}$ are bounded by $2$, and a rough upper bound for the degrees of $\binom{\mathcal{L}_{t}}{r}$ is $3r\binom{n}{r-1}=\mathcal{O}\left(n^{r-1}\right)$. We could then apply [Jan88, Theorem 2] to deduce asymptotic normality for the pattern count $X_{\pi}(\tau_{n})$, but we rather use Stein’s method to obtain a stronger result.
Proof of Theorem 2.14.
We aim to apply [Hof18, Theorem 3.5], which is based on the results of [CR10] and [Ros11], to the decomposition (1) of $X_{\pi}(\tau_{n})$. According to 7.1 and the discussion following it, $\binom{\mathcal{L}_{t}}{r}$ is a dependency graph for the family of random variables $\left(\mathbf{1}_{{\rm Perm}\left(Z_{i},i\in I\right)=\pi}\right)_{I\in\binom{\mathcal{V}_{t}}{r}}$. Thus we can use [Hof18, Theorem 3.5] with $B=1$, $D\leq 3r\binom{n}{r-1}$, and $N=\binom{n}{r}$. This yields the announced bound on the Kolmogorov distance.
@@
in probability, by Bienaymé–Chebychev’s inequality. If $\sigma>0$ then the previous bound on the Kolmogorov distance is a $\mathcal{O}(n^{-1/2})$. In particular it goes to 00 as $n\to\infty$, which implies convergence in distribution of the random variables. This concludes the proof. ∎
@@ 7.2 Computation of the variance
Fix a pattern $\pi\in\mathfrak{S}_{r}$. Throughout the proof we may omit the indices $\pi,p_{1},p_{2},n$. Thanks to Theorem 2.14, it suffices to compute the mean and variance of $X:=X_{\pi}(\tau)$. Let $\mathcal{P}=(Z_{i})_{i\in\mathcal{V}_{t}}$ be a geometric construction of $\tau_{n}$. The first idea is to approximate $X$ with a new variable
@@
Define $\mathcal{I}_{q}:=\left\{I\in\binom{\mathcal{V}_{t}}{r}\text{ indep., }\left|I\cap\mathcal{V}_{1}\right|=q\right\}$ and $c_{q}:=\left|\mathcal{I}_{q}\right|$. For any $I\in\mathcal{I}_{q}$, the variables $Z_{i},i\in I$ are independent by 7.1, and $\mathbf{P}\left({\rm Perm}\left(Z_{i},i\in I\right)=\pi\right)=\mu_{q}$. Therefore we can write $\mathbb{E}Y=\sum_{q=0}^{r}c_{q}\mu_{q}$. Since the number of non-independent $I\in\binom{\mathcal{V}_{t}}{r}$ is $\mathcal{O}(n^{r-1})$ and the total number of $I\in\binom{\mathcal{V}_{t}}{r}$ with $|I\cap\mathcal{V}_{1}|=q$ is
@@
we deduce $\mathbb{E}Y=\binom{n}{r}\mu+o(n^{r-1/2})$. It remains to study the variance of $Y$, and for this we rely on the method of $U$-statistics developed in [Hoe48]. The idea is to decompose $Y$ using its univariate projections. First define:
@@
When expanding $Y_{*}^{2}$, one can distinguish the pairs $(I,J)$ of $r$-sized independent subgraphs of $\mathcal{L}_{t}$ into three categories: the ones sharing no common vertex and connected by no edge; the ones sharing exactly one vertex or connected by exactly one edge; and the others. If $(I,J)$ is in the first category then $\mathbb{E}\left[\phi_{I}^{*}(Z_{i},i\in I)\phi_{J}^{*}(Z_{j},j\in J)\right]=0$ by independence (see 7.1). If $(I,J)$ is in the second category, write $(\alpha,\beta)\in I\times J$ for the unique pair of connected or equal vertices. Then $I\setminus\{\alpha\}$ is independent from $J\setminus\{\beta\}$ and we can write a.s.:
@@
$\left\{\begin{array}[]{ll}c_{i,q}=\binom{|\mathcal{V}_{1}|-1}{q-1}\binom{|\mathcal{V}_{t}\setminus\mathcal{V}_{1}|}{r-q}+\mathcal{O}(n^{r-2})=\binom{n}{r-1}\binom{r-1}{q-1}p_{1}^{q-1}(1\!-\!p_{1})^{r-q}+o(n^{r-3/2})\text{ uniformly in }i\in\mathcal{V}_{1}\vspace{.5em};\\ c_{i,q}=\binom{|\mathcal{V}_{1}|}{q}\binom{|\mathcal{V}_{t}\setminus\mathcal{V}_{1}|-1}{r-q-1}+\mathcal{O}(n^{r-2})=\binom{n}{r-1}\binom{r-1}{q}p_{1}^{q}(1\!-\!p_{1})^{r-q-1}+o(n^{r-3/2})\text{ uniformly in }i\in\mathcal{V}_{t}\setminus\mathcal{V}_{1}.\end{array}\right.$
@@
The previous calculation and (30) imply that $\left|\hat{Y}-\widetilde{Y}\right|=o(n^{r-1/2})$ a.s. for some deterministic $o$, hence as before we just need to compute the variance of $\widetilde{Y}$. Using 7.1 we can write:
@@
where $U,V,W$ are i.i.d. ${\rm Unif}\left([0,1]\right)$ variables. Thus $\mathrm{Var}\left[X\right]=\sigma^{2}n^{2r-1}+o(n^{2r-1})$ with $\sigma^{2}:=\Sigma_{\pi,\pi}^{p_{1},p_{2}}$ as defined in the statement of Theorem 2.15. Thanks to Theorem 2.14, this concludes the proof. ∎
Proof of 2.16.
Recall that conditionally given $t^{(n)}$, $\tau_{n}$ is a uniform $t^{(n)}$-cyclic permutation. Using Theorem 2.15, we can thus write for any $x\in\mathbb{R}$:
@@ 7.3 Non-degeneracy for involutive patterns
For any $p_{1}\in[0,1]$ and $\pi\in\mathfrak{S}_{r}$, the function $\psi_{\pi}^{p_{1}}$ is $(r\!-\!1)$-Lipschitz with respect to the $L^{1}$-norm on $[0,1]^{2}$. Moreover its restriction to $\{0\}\times[0,1]$ is a polynomial which is non-constant if $p_{1}<1$ and $r\geq 2$..
@@
Define $\mathcal{B}_{z,z^{\prime}}$ as the set of points in $[0,1]^{2}$ having x-coordinate between $u$ and $u^{\prime}$ or y-coordinate between $v$ and $v^{\prime}$. For the permutations ${\rm Perm}\left(\hat{Z}_{1},\dots,\hat{Z}_{r-1},z\right)$ and ${\rm Perm}\left(\hat{Z}_{1},\dots,\hat{Z}_{r-1},z^{\prime}\right)$ to differ, there needs to be a $\hat{Z}_{i}$ whose position with respect to $z$ (NW, NE, SW, or SE) is not the same as with respect to $z^{\prime}$. Therefore:
@@
and $\psi$ is indeed $(r\!-\!1)$-Lipschitz. Now fix $v\in[0,1]$ and let $Q$ be the number of $\hat{Z}_{i}^{\prime}s$, $i\leq r\!-\!1$ on the diagonal $\Delta$. The random variable $Q$ follows a ${\rm Binom}\left(r\!-\!1,p_{1}\right)$ law. Now let $q$, $j_{\rm SW}$, $j_{\rm SE}$, $j_{\rm NE}$, $j_{\rm NW}$, $k_{\rm SW}$, $k_{\rm NE}$ be nonnegative integers satisfying
@@
(the last line is redundant, but we keep it for clarity). Under this condition we write $A_{q,\bf j,k}(v)$ for the event that $Q=q$; $j_{\rm SW},j_{\rm SE},j_{\rm NE},j_{\rm NW}$ are the numbers of non-diagonal $\hat{Z}_{i}$’s in each quadrant defined by $(v,v)$; and $k_{\rm SW}$, $k_{\rm NE}$ are the numbers of diagonal $\hat{Z}_{i}$’s lying southwest, resp. northeast of $(v,v)$. If ${\rm Perm}\left(\hat{Z}_{1},\dots,\hat{Z}_{r-1},(0,v)\right)=\pi$ then $A_{q,\bf j,k}(v)$ is verified for some $q,\bf j,k$ satisfying (34). Moreover conditionally on $A_{q,\bf j,k}(v)$, the probability of this event does not depend on $v$. Indeed one can construct a straightforward coupling between $(\hat{Z}_{i})_{1\leq i\leq r-1}$ conditioned on $A_{q,\bf j,k}(v)$ and $(\hat{Z}_{i}^{\prime})_{1\leq i\leq r-1}$ conditioned on $A_{q,\bf j,k}(v^{\prime})$ such that a.s. ${\rm Perm}\left(\hat{Z}_{1},\dots,\hat{Z}_{r-1},(0,v)\right)={\rm Perm}\left(\hat{Z}_{1}^{\prime},\dots,\hat{Z}_{r-1}^{\prime},(0,v^{\prime})\right)$. Therefore:
@@
Conditionally given $V$, the variables $\psi(U,V)$ and $\psi(V,W)$ are independent, therefore the first term is null. Now let us study the second term. Since $\pi$ is an involution, for any $u_{1},v_{1},\dots,u_{r},v_{r}$ we have ${\rm Perm}\left((u_{1},v_{1}),\dots,(u_{r},v_{r})\right)=\pi$ if and only if ${\rm Perm}\left((v_{1},u_{1}),\dots,(v_{r},u_{r})\right)=\pi$. Since the distribution $p_{1}\mathrm{Leb}_{\Delta}+(1-p_{1})\mathrm{Leb}_{[0,1]^{2}}$ is symmetric with respect to $\Delta$, we deduce that $\psi(u,v)=\psi(v,u)$ for any $(u,v)\in[0,1]^{2}$. Hence:
@@
and in particular $(r-1)!^{2}\,\Sigma_{\pi,\pi}^{p_{1},p_{2}}\geq(1\!-\!p_{1}\!+\!p_{2})\mathrm{Var}\left[\psi(U,V)\right]$. By 7.2, $\psi$ is continuous and non-constant on $[0,1]^{2}$, therefore $\mathrm{Var}\left[\psi(U,V)\right]>0$. As $p_{1}<1$, this concludes the proof. ∎
@@ 7.4 Cycle types with few fixed points
where $\hat{Z}_{1},\dots,\hat{Z}_{r-1}$ are i.i.d. ${\rm Unif}\left([0,1]^{2}\right)$ variables, and the matrix of Theorem 2.15 rewrites as:
@@
where $U,V,W$ are i.i.d. ${\rm Unif}\left([0,1]\right)$ variables. To establish Theorem 2.18, we need to study the function $\psi$ a bit more.
@@
For any distinct $u_{1},v_{1},\dots,u_{r},v_{r}\in[0,1]$, we can write ${\rm Perm}\left((u_{1},v_{1}),\dots,(u_{r},v_{r})\right)=\sigma_{2}^{-1}\circ\sigma_{1}$ where $\sigma_{1},\sigma_{2}\in\mathfrak{S}_{r}$ are caracterized by $u_{\sigma_{1}(1)}<\dots<u_{\sigma_{1}(r)}$ and $v_{\sigma_{2}(1)}<\dots<v_{\sigma_{2}(r)}$. If $(U_{1},V_{1}),\dots,(U_{r},V_{r})$ are i.i.d. ${\rm Unif}\left([0,1]^{2}\right)$ variables then $\sigma_{1},\sigma_{2}$ are i.i.d. uniform permutations of size $r$. Conditionally given $(V_{1},\dots,V_{r})$, $\sigma_{2}$ is determined and $\sigma_{1}$ is still uniformly random in $\mathfrak{S}_{r}$. Subsequently
@@
[…]
For this event, say $A_{i}(u,v)$, to be realized, there needs to be exactly $i\!-\!1$ points left of $(u,v)$ and $\pi(i)\!-\!1$ below $(u,v)$. More precisely there needs to be exactly $j_{\mathrm{SW}},j_{\mathrm{SE}},j_{\mathrm{NE}},j_{\mathrm{NW}}$ points in each quadrant defined by $(u,v)$, where these quantities depend only on $\pi$ and $i$. Let $B_{i}(u,v)$ be this last event; its probability can be written as $\binom{r-1}{j_{\mathrm{SW}},\,j_{\mathrm{SE}},\,j_{\mathrm{NE}},\,j_{\mathrm{NW}}}u^{i-1}(1-u)^{r-i}v^{\pi(i)-1}(1-v)^{r-\pi(i)}$. Then conditionally on $B_{i}(u,v)$, the probability of $A_{i}(u,v)$ does not depend on $(u,v)$ anymore. Indeed by a coupling argument, one can e.g. see that $\mathbf{P}\left(A_{i}(u,v)\big{|}B_{i}(u,v)\right)$ is the probability that $j_{\mathrm{SW}}$ uniform points in $[0,1/2]^{2}$, $j_{\mathrm{SE}}$ uniform points in $[1/2,1]\!\times\![0,1/2]$, $j_{\mathrm{NE}}$ uniform points in $[1/2,1]^{2}$, and $j_{\mathrm{NW}}$ uniform points in $[0,1/2]\!\times\![1/2,1]$, all independent, form the permutation induced by $\pi$ on $[r]\setminus\{i\}$. Hence $\mathbf{P}\left(A_{i}(u,v)\big{|}B_{i}(u,v)\right)$ equals a positive constant, independent of $(u,v)$, and we deduce that
@@
Suppose that $\psi^{\rm sym}$ is constant on $[0,1]^{2}$. Then by setting $u=0$ we see that the polynomial $Q_{1}(v)$ is constant in $v$. If $\pi(1)>1$ then $\pi^{-1}(1)>1$ and the lowest degree term of $Q_{1}$ has positive degree and coefficient, which contradicts the fact that $Q_{1}$ is constant. Therefore $\pi(1)=1$. Likewise by setting $u=1$ we see that $Q_{r}$ is constant, which implies $\pi(r)=1$ or $\pi^{-1}(r)=1$. Hence $r=1$ and this concludes the proof. ∎
Proof of Theorem 2.18.
First, let us show non-degeneracy. Fix $p_{2}\in[0,1]$, $r\geq 2$, $\pi\in\mathfrak{S}_{r}$. Since $\psi_{\pi}$ is continuous and not constant on $[0,1]^{2}$ according to 7.4, it holds that $\mathrm{Var}\left[\psi_{\pi}(U,V)\right]>0$.
By Cauchy–Schwarz inequality, $\left|\mathrm{Cov}\left[\psi_{\pi}(U,V),\psi_{\pi}(V,U)\right]\right|\leq\mathrm{Var}\left[\psi_{\pi}(U,V)\right]$. Suppose by contradiction that $\Sigma_{\pi,\pi}^{p_{2}}=0$. Then necessarily $p_{2}=1$ and $\mathrm{Cov}\left[\psi_{\pi}(U,V),\psi_{\pi}(V,U)\right]=-\mathrm{Var}\left[\psi_{\pi}(U,V)\right]$. This equality case can only happen if $\psi_{\pi}(U,V)$ and $\psi_{\pi}(V,U)$ are linearly dependent, and here we would have $\psi_{\pi}(U,V)=-\psi_{\pi}(V,U)$ almost surely. However it follows from 7.4 that $\psi_{\pi}(U,V)+\psi_{\pi}(V,U)$ is not a.s. constant, and this proves by contradiction that $\Sigma_{\pi,\pi}^{p_{2}}>0$.
@@
Hence, up to a multiplicative constant, $\Sigma^{p_{2}}$ is the Gram matrix associated to the family $(\chi_{\pi})_{\pi\in\mathfrak{S}_{r}}$ in $L^{2}([0,1]^{2})$. Therefore the rank of $\Sigma^{p_{2}}$ equals the rank of this family. Recall the notation of 7.4. For any $(u,v)$ we have:

### v4

@@ 7.1 Dependency graphs and asymptotic normality
Let $t$ be a cycle type of size $n$, and $\mathcal{L}_{t}$ be the graph defined in Section˜1.3. The following lemma is a direct consequence of the definitions.
@@
In the geometric construction of ˜1.3, the graph $\mathcal{L}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$ is a dependency graph for the variables $(Z_{i})_{i\in\mathcal{V}_{t}}$.
Now we can use $\mathcal{L}_{t}$ to construct another dependency graph. Let $\binom{\mathcal{V}_{t}}{r}$ denote the set of subsets of size $r$ in $\mathcal{V}_{t}$. Define a graph $\binom{\mathcal{L}_{t}}{r}$ with vertices $\binom{\mathcal{V}_{t}}{r}$ as follows: $I,J\in\binom{\mathcal{V}_{t}}{r}$ share an edge if and only if there exist $i\in I$ and $j\in J$ such that $i=j$ or $i\sim j$ in $\mathcal{L}_{t}$. Using ˜7.1, it is straightforward to check that $\binom{\mathcal{L}_{t}}{r}$ is a dependency graph for the family of random variables $\left(\big.(Z_{i})_{i\in I}\right)_{I\in\binom{\mathcal{V}_{t}}{r}}$.
The idea that the variables $(Z_{i})_{i\in\mathcal{V}_{t}}$ are “partly dependent” is motivated by the fact that these graphs have low degrees: the degrees of $\mathcal{L}_{t}$ are bounded by $2$, and a rough upper bound for the degrees of $\binom{\mathcal{L}_{t}}{r}$ is $3r\binom{n}{r-1}=\mathcal{O}\left(n^{r-1}\right)$. We could then apply [J88, Theorem 2] to deduce asymptotic normality for the pattern count $X_{\pi}(\tau_{n})$, but we rather use Stein’s method to obtain a stronger result.
Proof of Theorem˜2.14.
We aim to apply [H18, Theorem 3.5], which is based on the results of [CR10] and [R11], to the decomposition (1) of $X_{\pi}(\tau_{n})$. According to ˜7.1 and the discussion following it, $\binom{\mathcal{L}_{t}}{r}$ is a dependency graph for the family of random variables $\left(\mathbf{1}_{{\rm Perm}\left(Z_{i},i\in I\right)=\pi}\right)_{I\in\binom{\mathcal{V}_{t}}{r}}$. Thus we can use [H18, Theorem 3.5] with $B=1$, $D\leq 3r\binom{n}{r-1}$, and $N=\binom{n}{r}$. This yields the announced bound on the Kolmogorov distance.
@@
in probability, by Bienaymé–Chebychev’s inequality. If $\sigma>0$ then the previous bound on the Kolmogorov distance is a $\mathcal{O}(n^{-1/2})$. In particular it goes to 0 as $n\to\infty$, which implies convergence in distribution of the random variables. This concludes the proof. ∎
@@ 7.2 Computation of the variance
Fix a pattern $\pi\in\mathfrak{S}_{r}$. Throughout the proof we may omit the indices $\pi,p_{1},p_{2},n$. Thanks to Theorem˜2.14, it suffices to compute the mean and variance of $X:=X_{\pi}(\tau)$. Let $\mathcal{P}=(Z_{i})_{i\in\mathcal{V}_{t}}$ be a geometric construction of $\tau_{n}$. The first idea is to approximate $X$ with a new variable
@@
Define $\mathcal{I}_{q}:=\left\{I\in\binom{\mathcal{V}_{t}}{r}\text{ indep., }\left|I\cap\mathcal{V}_{1}\right|=q\right\}$ and $c_{q}:=\left|\mathcal{I}_{q}\right|$. For any $I\in\mathcal{I}_{q}$, the variables $Z_{i},i\in I$ are independent by ˜7.1, and $\mathbf{P}\left({\rm Perm}\left(Z_{i},i\in I\right)=\pi\right)=\mu_{q}$. Therefore, we can write $\mathbb{E}Y=\sum_{q=0}^{r}c_{q}\mu_{q}$. Since the number of non-independent $I\in\binom{\mathcal{V}_{t}}{r}$ is $\mathcal{O}(n^{r-1})$ and the total number of $I\in\binom{\mathcal{V}_{t}}{r}$ with $|I\cap\mathcal{V}_{1}|=q$ is
@@
we deduce $\mathbb{E}Y=\binom{n}{r}\mu+o(n^{r-1/2})$. It remains to study the variance of $Y$, and for this we rely on the method of $U$-statistics developed in [H48]. The idea is to decompose $Y$ using its univariate projections. First define:
@@
When expanding $Y_{*}^{2}$, one can distinguish the pairs $(I,J)$ of $r$-sized independent subgraphs of $\mathcal{L}_{t}$ into three categories: the ones sharing no common vertex and connected by no edge; the ones sharing exactly one vertex or connected by exactly one edge; and the others. If $(I,J)$ is in the first category then $\mathbb{E}\left[\phi_{I}^{*}(Z_{i},i\in I)\phi_{J}^{*}(Z_{j},j\in J)\right]=0$ by independence (see ˜7.1). If $(I,J)$ is in the second category, write $(\alpha,\beta)\in I\times J$ for the unique pair of connected or equal vertices. Then $I\setminus\{\alpha\}$ is independent of $J\setminus\{\beta\}$ and we can write a.s.:
@@
$\left\{\begin{array}[]{ll}c_{i,q}=\binom{|\mathcal{V}_{1}|-1}{q-1}\binom{|\mathcal{V}_{t}\setminus\mathcal{V}_{1}|}{r-q}+\mathcal{O}(n^{r-2})=\binom{n}{r-1}\binom{r-1}{q-1}p_{1}^{q-1}(1\!-\!p_{1})^{r-q}+o(n^{r-3/2})\text{ uniformly in }i\in\mathcal{V}_{1}\vskip 5.0pt;\\ c_{i,q}=\binom{|\mathcal{V}_{1}|}{q}\binom{|\mathcal{V}_{t}\setminus\mathcal{V}_{1}|-1}{r-q-1}+\mathcal{O}(n^{r-2})=\binom{n}{r-1}\binom{r-1}{q}p_{1}^{q}(1\!-\!p_{1})^{r-q-1}+o(n^{r-3/2})\text{ uniformly in }i\in\mathcal{V}_{t}\setminus\mathcal{V}_{1}.\end{array}\right$.
@@
The previous calculation and (38) imply that $\left|\hat{Y}-\widetilde{Y}\right|=o(n^{r-1/2})$ a.s. for some deterministic $o$, hence as before we just need to compute the variance of $\widetilde{Y}$. Using ˜7.1 we can write:
@@
where $U,V,W$ are i.i.d. ${\rm Unif}\left([0,1]\right)$ variables. Thus $\mathrm{Var}\left[X\right]=\sigma^{2}n^{2r-1}+o(n^{2r-1})$ with $\sigma^{2}:=\Sigma_{\pi,\pi}^{p_{1},p_{2}}$ as defined in the statement of Theorem˜2.15. Thanks to Theorem˜2.14, this concludes the proof. ∎
Proof of ˜2.16.
Recall that conditionally given $t^{(n)}$, $\tau_{n}$ is a uniform $t^{(n)}$-cyclic permutation. Using Theorem˜2.15, we can thus write for any $x\in\mathbb{R}$:
@@ 7.3 Non-degeneracy for involutive patterns
For any $p_{1}\in[0,1]$ and $\pi\in\mathfrak{S}_{r}$, the function $\psi_{\pi}^{p_{1}}$ is (r1)(r\!-\!1)-Lipschitz with respect to the $L^{1}$-norm on $0,1]^{2}$. Moreover, its restriction to $\{0\}\times[0,1]$ is a polynomial which is non-constant if $p_{1}<1$ and $r\geq 2$..
@@
Define $\mathcal{B}_{z,z^{\prime}}$ as the set of points in $0,1]^{2}$ having x-coordinate between $u$ and $u^{\prime}$ or y-coordinate between $v$ and $v^{\prime}$. For the permutations ${\rm Perm}\left(\hat{Z}_{1},\dots,\hat{Z}_{r-1},z\right)$ and ${\rm Perm}\left(\hat{Z}_{1},\dots,\hat{Z}_{r-1},z^{\prime}\right)$ to differ, there needs to be a $\hat{Z}_{i}$ whose position with respect to $z$ (NW, NE, SW, or SE) is not the same as with respect to $z^{\prime}$. Therefore:
@@
and $\psi$ is indeed (r1)(r\!-\!1)-Lipschitz. Now fix $v\in[0,1]$ and let $Q$ be the number of $\hat{Z}_{i}^{\prime}s$, $i\leq r\!-\!1$ on the diagonal $\Delta$. The random variable $Q$ follows a ${\rm Binom}\left(r\!-\!1,p_{1}\right)$ law. Now let $q$, $j_{\rm SW}$, $j_{\rm SE}$, $j_{\rm NE}$, $j_{\rm NW}$, $k_{\rm SW}$, $k_{\rm NE}$ be nonnegative integers satisfying
@@
(the last line is redundant, but we keep it for clarity). Under this condition we write $A_{q,\bf j,k}(v)$ for the event that $Q=q$; $j_{\rm SW},j_{\rm SE},j_{\rm NE},j_{\rm NW}$ are the numbers of non-diagonal $\hat{Z}_{i}$’s in each quadrant defined by $(v,v)$; and $k_{\rm SW}$, $k_{\rm NE}$ are the numbers of diagonal $\hat{Z}_{i}$’s lying southwest, resp. northeast of $(v,v)$. If ${\rm Perm}\left(\hat{Z}_{1},\dots,\hat{Z}_{r-1},(0,v)\right)=\pi$ then $A_{q,\bf j,k}(v)$ is verified for some $q,\bf j,k$ satisfying (42). Moreover, conditionally on $A_{q,\bf j,k}(v)$, the probability of this event does not depend on $v$. Indeed one can construct a straightforward coupling between $(\hat{Z}_{i})_{1\leq i\leq r-1}$ conditioned on $A_{q,\bf j,k}(v)$ and $(\hat{Z}_{i}^{\prime})_{1\leq i\leq r-1}$ conditioned on $A_{q,\bf j,k}(v^{\prime})$ such that a.s. ${\rm Perm}\left(\hat{Z}_{1},\dots,\hat{Z}_{r-1},(0,v)\right)={\rm Perm}\left(\hat{Z}_{1}^{\prime},\dots,\hat{Z}_{r-1}^{\prime},(0,v^{\prime})\right)$. Therefore:
@@
Conditionally given $V$, the variables $\psi(U,V)$ and $\psi(V,W)$ are independent, and thus the first term is null. Now let us study the second term. Since $\pi$ is an involution, for any $u_{1},v_{1},\dots,u_{r},v_{r}$ we have ${\rm Perm}\left((u_{1},v_{1}),\dots,(u_{r},v_{r})\right)=\pi$ if and only if ${\rm Perm}\left((v_{1},u_{1}),\dots,(v_{r},u_{r})\right)=\pi$. Since the distribution $p_{1}\mathrm{Leb}_{\Delta}+(1-p_{1})\mathrm{Leb}_{[0,1]^{2}}$ is symmetric with respect to $\Delta$, we deduce that $\psi(u,v)=\psi(v,u)$ for any $(u,v)\in[0,1]^{2}$. Hence:
@@
and in particular $(r-1)!^{2}\,\Sigma_{\pi,\pi}^{p_{1},p_{2}}\geq(1\!-\!p_{1}\!+\!p_{2})\mathrm{Var}\left[\psi(U,V)\right]$. By ˜7.2, $\psi$ is continuous and non-constant on $0,1]^{2}$, and thus $\mathrm{Var}\left[\psi(U,V)\right]>0$. As $p_{1}<1$, this concludes the proof. ∎
@@ 7.4 Cycle types with few fixed points
where $\hat{Z}_{1},\dots,\hat{Z}_{r-1}$ are i.i.d. ${\rm Unif}\left([0,1]^{2}\right)$ variables, and the matrix of Theorem˜2.15 rewrites as:
@@
where $U,V,W$ are i.i.d. ${\rm Unif}\left([0,1]\right)$ variables. To establish Theorem˜2.18, we need to study the function $\psi$ a bit more.
@@
For any distinct $u_{1},v_{1},\dots,u_{r},v_{r}\in[0,1]$, we can write ${\rm Perm}\left((u_{1},v_{1}),\dots,(u_{r},v_{r})\right)=\sigma_{2}^{-1}\circ\sigma_{1}$ where $\sigma_{1},\sigma_{2}\in\mathfrak{S}_{r}$ are characterized by $u_{\sigma_{1}(1)}<\dots<u_{\sigma_{1}(r)}$ and $v_{\sigma_{2}(1)}<\dots<v_{\sigma_{2}(r)}$. If $(U_{1},V_{1}),\dots,(U_{r},V_{r})$ are i.i.d. ${\rm Unif}\left([0,1]^{2}\right)$ variables then $\sigma_{1},\sigma_{2}$ are i.i.d. uniform permutations of size $r$. Conditionally given $(V_{1},\dots,V_{r})$, $\sigma_{2}$ is determined and $\sigma_{1}$ is still uniformly random in $\mathfrak{S}_{r}$. Subsequently
@@
$\psi_{i}(u,v):=\mathbf{P}\left({\rm Perm}\left(\hat{Z}_{1},\dots,\hat{Z}_{r-1},(u,v)\right)=\pi\;,\;(u,v)\text{ is $i$-th from the left}\right).$
For this event, say $A_{i}(u,v)$, to be realized, there needs to be exactly i1i\!-\!1 points left of $(u,v)$ and $\pi(i)\!-\!1$ below $(u,v)$. More precisely there needs to be exactly $j_{\mathrm{SW}},j_{\mathrm{SE}},j_{\mathrm{NE}},j_{\mathrm{NW}}$ points in each quadrant defined by $(u,v)$, where these quantities depend only on $\pi$ and $i$. Let $B_{i}(u,v)$ be this last event; its probability can be written as $\binom{r-1}{j_{\mathrm{SW}},\,j_{\mathrm{SE}},\,j_{\mathrm{NE}},\,j_{\mathrm{NW}}}u^{i-1}(1-u)^{r-i}v^{\pi(i)-1}(1-v)^{r-\pi(i)}$. Then conditionally on $B_{i}(u,v)$, the probability of $A_{i}(u,v)$ does not depend on $(u,v)$ anymore. Indeed by a coupling argument, one can e.g. see that $\mathbf{P}\left(A_{i}(u,v)\big|B_{i}(u,v)\right)$ is the probability that $j_{\mathrm{SW}}$ uniform points in $0,1/2]^{2}$, $j_{\mathrm{SE}}$ uniform points in $1/2,1]\!\times\![0,1/2]$, $j_{\mathrm{NE}}$ uniform points in $1/2,1]^{2}$, and $j_{\mathrm{NW}}$ uniform points in $0,1/2]\!\times\![1/2,1]$, all independent, form the permutation induced by $\pi$ on $r]\setminus\{i\}$. Hence $\mathbf{P}\left(A_{i}(u,v)\big|B_{i}(u,v)\right)$ equals a positive constant, independent of $(u,v)$, and we deduce that
@@
Suppose that $\psi^{\rm sym}$ is constant on $0,1]^{2}$. Then by setting $u=0$ we see that the polynomial $Q_{1}(v)$ is constant in $v$. If $\pi(1)>1$ then $\pi^{-1}(1)>1$ and the lowest degree term of $Q_{1}$ has positive degree and coefficient, which contradicts the fact that $Q_{1}$ is constant. Hence $\pi(1)=1$. Likewise by setting $u=1$ we see that $Q_{r}$ is constant, which implies $\pi(r)=1$ or $\pi^{-1}(r)=1$. Hence $r=1$ and this concludes the proof. ∎
Proof of Theorem˜2.18.
First, let us show non-degeneracy. Fix $p_{2}\in[0,1]$, $r\geq 2$, $\pi\in\mathfrak{S}_{r}$. Since $\psi_{\pi}$ is continuous and not constant on $0,1]^{2}$ according to ˜7.4, it holds that $\mathrm{Var}\left[\psi_{\pi}(U,V)\right]>0$.
By Cauchy–Schwarz inequality, $\left|\mathrm{Cov}\left[\psi_{\pi}(U,V),\psi_{\pi}(V,U)\right]\right|\leq\mathrm{Var}\left[\psi_{\pi}(U,V)\right]$. Suppose by contradiction that $\Sigma_{\pi,\pi}^{p_{2}}=0$. Then necessarily $p_{2}=1$ and $\mathrm{Cov}\left[\psi_{\pi}(U,V),\psi_{\pi}(V,U)\right]=-\mathrm{Var}\left[\psi_{\pi}(U,V)\right]$. This equality case can only happen if $\psi_{\pi}(U,V)$ and $\psi_{\pi}(V,U)$ are linearly dependent, and here we would have $\psi_{\pi}(U,V)=-\psi_{\pi}(V,U)$ almost surely. However it follows from ˜7.4 that $\psi_{\pi}(U,V)+\psi_{\pi}(V,U)$ is not a.s. constant, and this proves by contradiction that $\Sigma_{\pi,\pi}^{p_{2}}>0$.
@@
Hence, up to a multiplicative constant, $\Sigma^{p_{2}}$ is the Gram matrix associated to the family $(\chi_{\pi})_{\pi\in\mathfrak{S}_{r}}$ in $L^{2}([0,1]^{2})$. Therefore, the rank of $\Sigma^{p_{2}}$ equals the rank of this family. Recall the notation of ˜7.4. For any $(u,v)$ we have: